What are E°



cell and ΔG° of a redox reaction at 25°C for



which n = 1 and K = 5.0×10−6?

1
Expert's answer
2021-11-19T09:24:02-0500

∆Gº = -RT ln K


R = 8.314 J/Kmol


T = 25 + 273 = 298K


K = 5x10-6


∆Gº = -(8.314)(298)(11.1)


∆Gº = - 27,500 J




∆Gº = -nFEº


∆Gº = -27.5 kJ


F = 96,485 C/mol e-


n = moles of e- = 1


Eºcell = -27500 / (-1)(96485) = 0.285 V